Two parallel wires A and B 10cm apart and carry currents of 10 A and 6 A respectively. Find the force on a 2.0 m length of wire B, if the currents are in the same direction

1
Expert's answer
2022-01-02T18:23:59-0500

"F=\frac{\mu_0I_1I_2l}{2\pi d},"

"F=\frac{2\cdot 10\cdot 6\cdot 2}{0.1}=2.4~kN."